Next year's earnings are estimated to be $5. The company plans to reinvest 25% of its earnings at 20%. If the cost of equity is 12%, what is the present value of growth opportunities?
Answer #1

Growth = Reinvestment * 20% = 25%*20% = 5%

Price = Dividend Next Year/( Cost of Equity - Growth) = 5*(1-25%)/(12%-5%) = 53.57

PVGO = Price - Earning/Cost of Equity = 53.57 - 5/12% = 11.90

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well
Answer #2

To calculate the present value of growth opportunities (PVGO), we need to determine the value of the reinvested earnings. PVGO represents the additional value generated by the company's future investments and growth prospects.

The formula to calculate PVGO is as follows:

PVGO = (Reinvested Earnings) / Cost of Equity

Given:

  • Estimated earnings next year: $5

  • Reinvestment rate: 25% (0.25)

  • Reinvestment return: 20% (0.20)

  • Cost of equity: 12% (0.12)

Step 1: Calculate the amount of earnings to be reinvested: Reinvested Earnings = Estimated earnings next year * Reinvestment rate

Reinvested Earnings = $5 * 0.25 = $1.25

Step 2: Calculate the present value of the reinvested earnings: PVGO = Reinvested Earnings / Cost of Equity

PVGO = $1.25 / 0.12

PVGO ≈ $10.42

Therefore, the present value of growth opportunities (PVGO) is approximately $10.42. This value represents the additional value generated by the company's future investments, considering the reinvestment rate, reinvestment return, and cost of equity.


answered by: mervetokaz